OpenStudy (anonymous):

Can someone explain me how to solve this? Given e^x y+y^2 x - sinx=6 Find dy/dx

OpenStudy (anonymous):

go to tutor.com they can help you

OpenStudy (anonymous):

implicit diff think of \[y=f(x),y'=f'(x)\] and use the product, quotient and chain rule where applicable.

OpenStudy (anonymous):

\[y+y^2x-\sin(x)=6\] \[y'+2yy'x+y^2-\cos(x)=0\] in this case, the solve for y' with some algebra

OpenStudy (anonymous):

your way smarter than me lol but im only in 8th grade

OpenStudy (anonymous):

you get \[y'+2yy'x=\cos(x)-y^2\] \[y'=\frac{\cos(x)-y^2}{1+2xy}\]

OpenStudy (anonymous):

thank you so much u saved my life!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

e^x was in the original problem, but you didn't have it in your answer